### v2

# New Tool to Detect Inhomogeneous Chiral Symmetry Breaking

Abstract
In this letter, we discuss a novel method to search for inhomogeneous chiral symmetry breaking in theories with fermions. The prime application we have in mind is QCD, but the method is also applicable for other theories, including solid state applications. It is based on an extension of the chiral susceptibility to inhomogeneous phases and it works as a stability analysis. Our method allows us to determine when homogeneous solutions have the tendency to create spatially modulated condensates. As proof of principle, we apply this technique to a rainbow-ladder QCD model and confirm the conjectured shape of the instability region published in Ref. [1].
I Introduction
Chiral Symmetry Breaking (SB) is one of the most consequential phenomena associated with strong interactions. With massless quarks, the QCD Lagrangian is manifestly chirally symmetric, however the ground state breaks chiral symmetry. For two massless quark flavours, at high temperatures the system undergoes a second-order phase transition to the symmetric phase, which becomes a cross-over at finite quark masses [2, 3]. What happens to chiral symmetry at large chemical potentials, however, is less clear. One possibility is that the system undergoes a transition to a phase where chiral symmetry is broken inhomogeneously, that is, the order parameter is a periodic function of space. These phases have been seen in models of strong interactions [4, 5, 6, 7]. If real, these phases might occur in a region of the phase diagram that is connected to the physics of neutron stars and neutron-star-mergers as well as with low-energy heavy-ion collisions. In fact, it has already been argued [8, 9, 10, 11] that moat regimes (a broader but correlated phenomenon to inhomogeneous phases, see also [12, 13]) would leave experimental signatures in scattering experiments.
From a theoretical perspective, it is highly non-trivial to obtain solid information on the existence of inhomogeneous phases at large chemical potential. Lattice calculations of QCD are notoriously difficult if not prohibited in this region because of the sign problem [14]. Therefore, some attention has been paid to simpler (and partly lower-dimensional) models, such as Gross-Neveu (GN) or Nambu-Jona-Lasinio (NJL) type models [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], which are accessible to both, continuum and lattice methods. Although much has been learned from the models, in the end one needs to study the problem within QCD.
In general, two strategies have been employed so far. Using a specific ansatz for the shape of the inhomogeneous phase, one can calculate the free energy in this configuration and compare with the homogeneous one. One of the most commonly used ansatzes, e.g., is the so-called Chiral Density Wave (CDW) [5]. Here, the condensates oscillate with spatial variable
where is the third Pauli matrix in flavour space and is the wave vector of the modulation. In the NJL-model the amplitude of this modulation depends on the coupling constant to the scalar and pseudo-scalar sectors [5]. Alternatively, one can perform what is called a stability analysis which probes the stability of the homogeneous phase with respect to very small inhomogeneous perturbations. This technique has the benefit that it is agnostic with respect to the shape of the inhomogeneous phase. On the other hand, since it is based on small perturbations, it could miss instabilities with respect to large fluctuations. Nevertheless, these two techniques typically agree on second-order phase boundaries.
Both, the ansatz and the stability analysis techniques face severe additional difficulties in QCD compared to their application in simple models. Due to the non-locality of the self-energy the expectation values of quark bilinears are functions of two momentum variables in an inhomogeneous phase. To our knowledge, Ref. [30] is the only attempt so far to deal with these difficulties in a functional approach to QCD via Dyson-Schwinger Equations (DSEs). Using an ansatz for the quark propagator which yields condensates like Eq. (1), they show that spatially inhomogeneous solutions to their set of (truncated) DSEs appear at high chemical potentials.
The stability analysis framework has only recently been generalised from model applications to QCD [31, 1]. Although agnostic in general, our specific formulation of the framework made necessary by the added layer of complexity of QCD relied on the specific shape of the perturbations (what we called the “test-function”). Furthermore, we were only able to apply the analysis to a very limited region of the phase diagram, namely the left spinodal line. The details are explained in Refs. [31, 1], here we just note that the shape of our test-functions has constraints, which we can only exactly enforce in this region. These shortcomings notwithstanding, the general methodology of stability analysis is still highly promising and its first application to a simple truncation of QCD [1] shows the tendency to create inhomogeneous condensates.
In this letter, we overcome these shortcomings by introducing a novel technique to perform a stability analysis. This new method is applicable to any theory, including simple models, arbitrary truncations of QCD but also effective theories in solid state physics. The technique does not depend on any test-functions, and it is not restricted to any region of the phase diagram. It is based on a well-known method to identify the second-order homogeneous phase boundaries, namely calculating the chiral susceptibility (see e.g. Refs. [32, 33]). However, by performing a chiral rotation to the quark fields, we are able to compute an inhomogeneous chiral susceptibility and determine when the symmetric solution is unstable with respect to inhomogeneous chiral symmetry breaking.
The outline of the Letter is as follows. In the next section we will introduce this idea in the context of NJL. We show how it reproduces previous results. We then move to generalise it to QCD and show first proof-of-principle results in a simple truncation. Discussions and conclusions follow.
II Inhomogeneous Chiral Susceptibility in NJL
Let us illustrate the idea in the NJL model. Take the following standard Lagrangian (note that we will use the same Euclidian-space conventions as in Ref. [31] and as for the numerical values for the couplings, cutoff, and regularisation scheme, we will take the setup of Ref. [4])
We will work in the chiral limit of zero bare mass . In mean-field NJL, the inverse dressed fermion propagator
where is the dynamically generated mass, which depends on the chiral condensate, and is an order parameter for SB. One way to resolve the phase diagram of the theory is to calculate the derivative of this order parameter with respect to the bare quark mass. For the symmetric phase, this is simply
where we denote the combination of the Matsubara sum of the frequencies and 3-momentum integration as . The susceptibility is negative where the symmetric phase is unstable with respect to small homogeneous breakings of chiral symmetry, and it will be positive otherwise. Close to a second-order chiral transition, the sign change of exactly finds the phase transition. Around a first-order transition, however, there is a region where both the restored and the broken solution are (meta-)stable. This is known as the spinodal region and its edges are known as the spinodal lines. In Figure 1, we show how the sign of neatly resolves the second-order transition and left-spinodal lines in NJL, i.e., the regions where the restored solution changes from unstable to stable.
Take now the following (chiral-rotated) modified quark fields [34]
The NJL-Lagrangian, Eq. (2), then transforms to
For non-zero , assuming a homogeneous condensate and , it is easy to show that we obtain a modulated just like Eq. (1),
Note that, when we omit the spatial arguments in the condensate, we mean that it is constant in space.
Under this transformation, the inverse fermion propagator for non-zero becomes
This again can be inverted and an inhomogeneous chiral susceptibility can be calculated to be the following
This expression is general, however, in practice we take the wave-vector taken to be fixed in the direction, i.e., . In Fig. 2 we show how this resolves exactly the phase boundary of the inhomogeneous phase as well: By varying we manage to find where the symmetric solution is unstable with respect to inhomogeneous SB, in regions where it is not unstable with respect to homogeneous SB. The convolution of the red regions in Fig. 2, that is, regions of negative for some value of , reproduces exactly the unstable region in Ref. [4].
Let us now proceed to extending this idea to truncations of QCD.
III QCD
As it was the case for the NJL example, first we need to find the chiral symmetric solution, in order to calculate its inhomogeneous chiral susceptibility. We take a rainbow-ladder model of QCD defined by the following quark Dyson-Schwinger Equation
Here, the bare line denotes the (inverse) bare propagator, whereas the line with the black dot represents the dressed (inverse) propagator including quantum fluctuations,
where . The self-interaction diagram contains a dressed gluon propagator (wiggly line) and two bare quark-gluon vertices. The effect of dressing the gluon propagator with momentum and one of the quark gluon vertices, , is modelled by two analytic functions and
Here, factors of the QCD coupling and renormalisation functions of the vertex and the quark have been absorbed into the effective couplings and . The transverse () and longitudinal () projectors are given by
As an example for such an effective running coupling we take the Qin-Chang model [35] with a thermal mass for the gluon calculated via hard-thermal and hard-dense loop techniques
This model has been applied to aspects of the QCD phase diagram before, see e.g. the review Ref. [36] for a summary.
As a first step in the analysis, we solve the DSE for the homogeneous restored phase, for which is exactly zero, all over the phase diagram. This is done in the standard formulation of QCD without the chiral rotation Eq. (5). We then force the appearance of a modulation by performing the rotation, which adds a term to the inverse propagator, i.e., in the QCD Lagrangian
where here refers to the covariant derivative. In the propagator, this extension induces several new terms which can be worked out exactly. Their details do not matter here. Each of these new terms is accompanied by its own dressing function. In an ansatz-based approach, one would have to go through the complicated and extremely costly exercise to solve the resulting set of DSEs self-consistently to get the stable CDW solution.
Is is, however, much simpler and numerically cheaper to determine the inhomogeneous chiral susceptibility in this framework. A trivial extrapolation of the calculation above will show that the inhomogeneous chiral susceptibility of the chiral solution is the following
Again, a non-trivial modulated chiral condensate is indicated if this function goes negative at finite . Comparing Eq. (17) with the corresponding one for the NJL-model, Eq. (9), we find the slight complication that the susceptibility now depends on the quark momentum and appears on both sides, and the equation has to be solved self-consistently. Note, however, that the quark dressing functions and on the right hand side are the ones calculated without the extra term present in the theory. This is a subtle and important point. We are not seeking for a self-consistent solution of the DSE with the extra term present. Instead, the question we want answered is: once we add , will this induce the formation of a Dirac scalar term to the quark propagator? If so, then the susceptibility diverges at the second-order phase boundary and an inhomogeneous phase will be formed.
In Fig. 3 we show the line where the susceptibility diverges for different values of . This analysis not only confirms our earlier findings from Ref. [1] but also significantly advances our results by moving beyond test-function-based stability analyses. Here, we conclusively show that the chiral solution is unstable to inhomogeneous perturbations, with a proto-Lifschitz point 111A proto-Lifschitz point is a term we define in Ref. [1] and it simply refers to the point where the stability boundary of the symmetric phase w.r.t. inhomogeneous fluctuations meets the left spinodal of the homogeneous first-order transition appearing near MeV. As seen in Fig. 3, we can definitively delineate the region bounded by the spinodal line on one side and the inhomogeneous convolution (black dotted) line on the other as a region where the chiral solution really is unstable against crystallisation. Importantly, we establish that for temperatures as low as 70 MeV, the instability region does not extend into the first-order transition. Since in this region the restored solution is not energetically favoured over the homogeneous broken solution, we identify this as a feature of a disfavored phase rather than an indicator of an emergent inhomogeneous phase. Note, however, that with this method we now can perform the stability analysis where the homogeneous restored solution is preferred over the homogeneous broken one, and, there, we identify that it is stable w.r.t. small inhomgeneities.
Exploring lower temperatures may reveal points where the instability line crosses the first-order phase transition. However, our current methodology—though substantially more precise and conclusive than prior approaches—becomes increasingly resource-intensive at lower temperatures since it requires solving the homogeneous DSE with large numerical precision. Thus, rather than extending this model analysis further, we prioritize more realistic truncations as a direction for future work, where our methodology can provide more rigorous insights. This work establishes a solid foundation and methodology to probe these complexities in the future, with higher-fidelity models.
IV Conclusions
In this work, we proposed a new method to analyse the stability of a chirally symmetric phase against the formation of an inhomogeneous condensate via chiral symmetry breaking. The method is based on an extension of the concept of chiral susceptibility toward probing the reaction of the theory to the presence of spatially modulated quark fields. As a proof of principle, we showed that our method quantitatively reproduces previous NJL-model results [4] on the presence of inhomogeneous phases and we confirmed previous expectations for instability in a rainbow-ladder model of QCD [1].
We would like to stress again, that this methodology to study inhomogeneous chiral symmetry breaking is very general. It does not depend on a test-function and it is equally applicable to any truncation of QCD, with no computational costs coming from complex truncations. Since it is not necessary to know the full momentum dependence of , we expect this should also be easily calculable with the functional renormalisation group.
Some studies suggest that additional quantum fluctuations could destabilise inhomogeneous condensates (see Ref. [38]). In that case, we expect the homogeneous solution to the DSE to capture the energetically preferred, stable configuration. Other studies indicate that the boundary between inhomogeneous and restored phases may be first-order (see Ref. [39]), limiting our ability to precisely determine the transition line. Nevertheless, our approach should still identify the corresponding spinodal.
One should, however, keep in mind that a stability analysis in general does not say much about the fully self-consistent solution of the system. It just shows that there is the tendency to create a mass term (break chiral symmetry) once we introduce a modulation, in this instance a CDW-like condensation. The actual shape of the modulation would eventually to be determined by a self-consistent solution of the DSEs for all dressing functions of the quark. This is a tremendous effort, but one that should be undertaken once our method does show instabilities in more advanced truncations to QCD than used in this work.
Finally, it is important to realise that such instabilities (or even their associated phenomenon, the moat regimes) are highly relevant as they might leave signatures in heavy-ion collisions [8]. Even if these phases turn out not to be exactly stable in full QCD, these signatures might still be present due to long-lived meta-stable realisations of these phases as the system moves towards equilibrium [40].
Acknowledgements
It is our pleasure to thank Fabian Rennecke, Julian Bernhardt, Ricardo Costa de Almeida, Renan Hirayama, and Curtis Abell for helpful discussions about the content of this manuscript. This work has been supported by the Alexander von Humboldt Foundation, the Deutsche Forschungsgemeinschaft (DFG) through the Collaborative Research Center TransRegio CRC-TR 211 “Strong-interaction matter under extreme conditions” and the individual grant FI 970/16-1.